The force of interest at any time t, (measured in years) is given by

 0.07−0.005t, 0≤t<5 0.06 − 0.003t, 5 ≤ t < 10

δ(t) =

What is the total accumulated in value at any time t (> 0) of investments of $100 at

0.03, t ≥ 10 times 0, 4 and 6?


1
Expert's answer
2022-01-28T06:09:08-0500

"a)100 \times A(0,5) \times A(5,10) \times A(10,15)
\\A=100\times e^{\int_0^5(0.07-0.005t)dt} \times e^{\int_5^{10}(0.06-0.003t)dt} \times e^{0.03 \times 5}\\
=100 \times e^{0.07t-\frac{0.005t^2}{2}|_0^5} \times e^{0.06t-\frac{0.003t^2}{2}|_5^{10}} \times e^{0.15}\\


=100 \times e^{0.35-0.0625} \times e^{0.45-0.2925} \times e^{0.15}\\
=100 \times e^{0.2875} \times e^{0.1575} \times e^{0.15}\\
=100 \times 1.3331 \times 1.1706 \times 1.1618\\
=\$ 181.30\\
at~t=0\\
A=100 \times e^{0.07 \times 5} \times e^{0.06 \times 5} \times e^{0.073\times 5}\\
A=100 \times 1.4191 \times 1.3499 \times 1.1618\\
A=\$ 222.56\\
at~t=4\\
=100\times e^{\int_4^5(0.07-0.005t)dt} \times e^{\int_5^{10}(0.06-0.003t)dt} \times e^{0.03 \times 5}\\
=100 \times e^{0.07t-\frac{0.005t^2}{2}|_4^5} \times e^{0.06t-\frac{0.003t^2}{2}|_5^{10}} \times e^{0.15}\\
=100 \times e^{0.2875-0.24} \times e^{0.45-0.2925} \times e^{0.15}\\
=100 \times e^{0.0475} \times e^{0.1575} \times e^{0.15}\\
=100 \times 1.0486 \times 1.1706 \times 1.1618\\
=\$ 142.61\\
at~t=6\\
=100 \times e^{\int_6^{10}(0.06-0.003t)dt} \times e^{0.03 \times 5}\\
=100 \times e^{0.06t-\frac{0.003t^2}{2}|_6^{10}} \times e^{0.15}\\
=100 \times e^{0.45-0.27} \times e^{0.15}\\
=100 \times e^{0.18} \times e^{0.15}\\
=100 \times 1.1972 \times 1.1618\\
=\$ 139.09\\"